# Pricing energy spread options with variance gamma-driven Ornstein-Uhlenbeck dynamics

### v2

@@ Abstract
We consider the pricing of energy spread options for spot prices following an exponential Ornstein-Uhlenbeck process driven by a sum of independent multivariate variance gamma processes, which gives rise to mean-reverting, infinite activity price dynamics. Within this class of driving processes, the Esscher transform is used to obtain an equivalent martingale measure with a focus on the weak variance alpha-gamma process. By deriving an analytic expression for the cumulant generating function of the innovation term, we obtain a pricing formula for forwards and apply the FFT method of Hurd and Zhou to price spread options. Lastly, we demonstrate how the model should be both estimated on energy prices under the real world measure and calibrated on forward or call prices, and provide numerical results for the pricing of spread options.
@@ 1 Introduction
with strike price $K$ and maturity time $T$. Energy prices tend to be highly correlated across commodities and regions. For example, electricity prices are related to the price of gas used to generate it. An option on this difference is called a spark spread, which are used by generators to hedge against unanticipated spikes in production costs. As another example, the physical interconnectors between the five regions of the National Electricity Market in Australia also induce a correlation, and an option on such a difference would be a regional spread. The survey by Carmona and Durrleman cardur03 catalogs the variety of spread options. In general, there is no known closed-form pricing formula for spread options, except when $K=0$ in a lognormal model where it reduces to the well-known Margrabe formula.
@@
where $\bf Z$} is the background driving Lévy process (BDLP), $\lambda>0$ is the speed of mean reversion, and the independent random vector ${\bf X}_{0}$ is the initial value. The Ornstein-Uhlenbeck part ensures mean reversion, while the Lévy part allows for jumps, and indeed the use of LDOUPs is widespread in energy price modeling. See for example BKMB ; benthbook ; BenthPir18 ; cf2005 ; sab2020 ; sab2023 .
The BDLP $\bf Z$} we use herein is a sum of independent multivariate variance gamma (VG) processes with a focus on the weak variance alpha-gamma (WVAG) process by Buchmann, Lu and Madan BLM17a as a subclass. A multivariate VG process is a multivariate Brownian motion $\bf B$} subordinated, or time-changed, by a univariate gamma subordinator, and such processes originated with Madan and Seneta MaSe90 and have been studied in BKMS16 ; FS07 . While a sum of VG processes does not necessarily preserve such a time-change interpretation, the WVAG process being originally defined via weak subordination does. It turns out to be equivalent to a process $\bf Z$} characterized in the following way: consider a subordinator $\bf T$} which is a sum of $G_{k}$ representing the idiosyncratic time-change for the $k$th price, and $\boldsymbol{\alpha}G_{0}$ representing a common time change for all prices, where $G_{0},\dots,G_{n}$ are independent gamma subordinators and $\boldsymbol{\alpha}\in[0,\infty)^{n}$, then when ${\bf T}(t)$ jumps by $\Delta{\bf T}(t)$, ${\bf Z}(t)$ jumps by the componentwise composition ${\bf B}(\Delta{\bf T}(t))$ in distribution, rather than pathwisely for traditional subordination. Consequently, the WVAG process, using weak subordination, preserves the essence of traditional subordination in a distributional or weak sense.
This provides a naturally multivariate process with a flexible dependence structure where the Brownian motion has arbitrary covariance in contrast to other approaches, such as using traditional subordination restricted to independent Brownian motions to remain a Lévy process as considered by Luciano and Semeraro LS10 , or by summing independent univariate VG processes $V_{0},\dots,V_{n}$ in the form ${\bf Z}=(\alpha_{1}V_{0}+V_{1},\dots,\alpha_{n}V_{0}+V_{n})$, an approach found in Ballotta and Bonfiglioni BB14 , Caldana and Fusai CaFu2013 , and Hurd and Zhou HZ10 , with the latter two in the context of spread option pricing. Our BLDP $\bf Z$} includes these cases. A similar way of summing univariate LDOUPs to produce a bivariate energy price for spread option pricing is also considered in (benthbook, , Section 9.2.1). Our LDOUP $\bf X$} include certain restrictions of this model, although both models are more general than each other in different ways.
@@
For the pricing of energy derivatives, we use the Esscher transform gs94 to obtain an equivalent martingale measure (EMM) or Esscher measure $\mathbb{Q}_{{\bf h}}$ indexed by the market price of risk (MPR) $\bf h$}. This is a common approach for Lévy-based models, in which the market is incomplete due to the presence of jumps. Energy prices such as electricity are nontraded assets in the sense that term is used in standard arbitrage theory since electricity bought or sold on the market is instantly consumed and cannot be held to construct a replicating portfolio for the valuation of derivatives. Consequently, it is not assumed that the discounted energy price process is a $\mathbb{Q}_{{\bf h}}$-martingale, a condition in an equity market that would ordinarily determine a unique $\bf h$}, and hence an EMM $\mathbb{Q}_{{\bf h}}$. Instead, in an energy market, $\bf h$} must be obtained by calibration, such as to observed forward or call prices.
@@
Here, we show how the BDLP $\bf Z$} changes under the Esscher transform into a sum of VG processes with modified parameters, which leads to the result that the class of WVAG processes is not closed under Esscher transform. In other related works, Buchmann et al. BKMS16 studies Esscher transforms for a class of strongly subordinated Brownian motions, and Michaelsen and Szimayer MiSz17 provides numerical results about pricing options on equities that follow a exponential WVAG process using the Esscher transform, although these rely on results that do not extend to energy derivatives. Our results are more general.
@@
We derive an analytical formula for the cumulant generating function (cgf) of the innovation term of a LDOUP where the BLDP is a multivariate VG process in terms of the dilogarithm function, which generalizes a result in the univariate case by Sabino sab2020 . This is a key result given in Theorem 3.3, and is surprising in light of Remark 3.6. Consequently, we obtain an analytical formula for the cgf of the log return along with conditions for this to be finite and when it can be analytically continued into the complex plane. We use this cgf to obtain a forward pricing formula, and combined with the Fourier-based methods of Carr and Madan CM99 and Hurd and Zhou HZ10 , it also gives pricing formulas for call and spread options, respectively. This approach to pricing spread options differs from Benth and Šaltytė-Benth bsb06 where the spread is modelled directly as a univariate LDOUP as opposed to being based on (1.1). The Hurd-Zhou method has also been studied for spread options on equities and interest rates in Alfeus and Schlögl AlSc2018 .
For the WVAG BDLP, we provide numerical results under both the true model and the fitted model. Under the true model, the spread option price is computed via the analytical formula for the cgf of the log return with the FFT method, and is compared to the Monte Carlo method, showing close agreement. We also provide an end-to-end implementation of the spread option pricing in a simulation study by fitting the model. Our model allows the seasonality and LDOUP parameters to be estimated under the real world measure $\mathbb{P}$, specifically using the MLE method by Lu Lu21 combined with the analytical formula for the cgf of the innovation term, while the MPR $\bf h$} is estimated under the Esscher measure $\mathbb{Q}_{{\bf h}}$, in the sense of calibrating the model to observed forward and call prices. Thus, we can coherently leverage both data on the $\mathbb{P}$-dynamics of the underlying and its derivative prices for model fitting, in contrast to the practice of fully calibrating the model under the risk-neutral measure for pricing energy derivatives gs22 ; sab2020 , and more commonly equity derivatives MCC98 ; sco . Finally, the fitted model can be used to price spread options over a panel of strikes. We find in our example that calibrating to calls is better than calibrating to forwards, and demonstrate the extent to which errors in estimating the model parameters propagate to pricing errors.
@@
The paper is organized as follows. Section 2 reviews the preliminaries, including notation, definitions of the VG and WVAG processes, and useful properties of LDOUPs. Section 3 presents the main probabilistic result, giving the analytical formula for the cgf of the innovation term of a LDOUP where the BDLP is multivariate VG. Section 4 focuses on the pricing of energy derivatives and is the main part of the paper, where we specify the model, show how the BDLP changes under the Esscher transform, obtain the analytical cgf of the log return, and finally apply this to derive pricing formulas for forwards, calls, and spread options. Section 5 details the implementation of our model and the pricing formulas, including simulation, estimation, calibration, and the pricing of spread options using the FFT method. Section 6 presents numerical results based on both the true parameters and a simulation study using fitted parameters. Section 7 concludes with some remarks on possible extensions.
@@ 2 Preliminaries
Let $\kappa_{{\bf X}}^{\mathbb{Q}}$, $\Phi_{{\bf X}}^{\mathbb{Q}}$ and $\Psi_{{\bf X}}^{\mathbb{Q}}$ denote the respective functions under $\mathbb{Q}$. Let ${\bf B}\sim BM^{n}(\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ be an $n$-dimensional Brownian motion with drift $\boldsymbol{\mu}\in\mathbb{R}^{n}$ and covariance ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$. For ${\bf t}=(t_{1},\dots,t_{n})\in[0,\infty)^{n}$, define
@@ 2.1 Variance gamma and related process
A VG convolution (VGC) or sum of VG processes ${\bf Z}\sim{VGC}^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ is defined by
${\bf Z}\stackrel{{\scriptstyle D}}{{=}}\boldsymbol{\eta}I+{\bf V}_{1}+\dots+{\bf V}_{m},$ (2.4)
@@
${\bf Z}\stackrel{{\scriptstyle D}}{{=}}\boldsymbol{\eta}I+{\bf V}_{0}+(V_{1},\dots,V_{n}),$ (2.5)
@@
which in turn allows the process to be simulated. The marginal components of ${\bf Z}=(Z_{1},\dots,Z_{n})$ are $Z_{k}\sim VG^{1}(1/\alpha_{k},\mu_{k},\allowbreak\Sigma_{kk},\eta_{k})$, which is a general univariate VG processes
@@ 2.2 Lévy-Driven Ornstein-Uhlenbeck Processes
The solution to the SDE (1.2) is
${\bf X}(t)=e^{-\lambda t}{\bf X}(s)+e^{-\lambda t}\int_{s}^{t}e^{\lambda u}\,{\rm d}{\bf Z}(\lambda u),\quad 0\leq s\leq t.$ (2.9)
For $0\leq s\leq t$, define the random vectors
${\bf Z}^{*}(s,t):=\int_{s}^{t}e^{\lambda u}\,{\rm d}{\bf Z}(\lambda u),\quad{\bf Z}^{*}(t):={\bf Z}^{*}(0,t).$
This notation is from vstt . Note that
@@
where $b=e^{-\lambda\Delta}$, and ${\bf Z}_{b}^{(k)}\stackrel{{\scriptstyle D}}{{=}}e^{-\lambda\Delta}{\bf Z}^{*}(\Delta)$ are the iid innovations. For convenience, we instead call ${\bf Z}^{*}(\Delta)$ the innovation term, and it has characteristic exponent is
@@
(see (saya, , Theorem 2.2)). This implies that if ${\bf Z}=\boldsymbol{\eta}I+{\bf Z}_{1}+\dots+{\bf Z}_{m}$, where ${\bf Z}_{1},\dots,{\bf Z}_{m}$ are independent Lévy processes, then

### v3

@@ Abstract
We consider the pricing of energy spread options for spot prices following an exponential Ornstein-Uhlenbeck process driven by a sum of independent multivariate variance gamma processes, which gives rise to mean-reverting, infinite activity price dynamics. Within this class of driving processes, the Esscher transform is used to obtain an equivalent martingale measure with a focus on the weak variance alpha-gamma process. By deriving an analytical formula for the cumulant generating function of the innovation term, we obtain a pricing formula for forwards and apply the FFT method of Hurd and Zhou to price spread options. Lastly, we demonstrate how the model should be both estimated on energy prices under the real world measure and calibrated on forward or call prices, and provide numerical results for the pricing of spread options.
@@ 1 Introduction
with strike price $K$ and maturity time $T$. Energy prices tend to be highly correlated across commodities and regions. For example, electricity prices are related to the price of gas used to generate it. An option on this difference is called a spark spread, which are used by generators to hedge against unanticipated spikes in production costs. As another example, the physical interconnectors between the five states of the National Electricity Market (NEM) in Australia also induce a correlation, and an option on such a difference would be a regional spread. The survey by Carmona and Durrleman cardur03 catalogs the variety of spread options. In general, there is no known closed-form pricing formula for spread options, except when $K=0$ in a lognormal model where it reduces to the well-known Margrabe formula.
@@
where $\bf Z$} is the background driving Lévy process (BDLP), $\lambda>0$ is the speed of mean reversion, and the independent random vector ${\bf X}_{0}$ is the initial value. The Ornstein-Uhlenbeck part ensures mean reversion, while the Lévy part allows for jumps, and indeed the use of LDOUPs is widespread in energy price modeling. For example, see BKMB ; benthbook ; BenthPir18 ; cf05 ; sab2020 ; sab2023 .
The BDLP $\bf Z$} we use herein is a sum of independent multivariate variance gamma (VG) processes with a focus on the weak variance alpha-gamma (WVAG) process by Buchmann, Lu and Madan BLM17a as a subclass. A multivariate VG process is a multivariate Brownian motion $\bf B$} subordinated, or time-changed, by a univariate gamma subordinator $G$, and such processes originated with Madan and Seneta MaSe90 and have been studied in BKMS16 ; FS07 . While a sum of VG processes does not necessarily preserve such a time-change interpretation, the WVAG process being originally defined via weak subordination does. It turns out to be equivalent to a process $\bf Z$} characterized in the following way: consider a subordinator $\bf T$} which is a sum of $G_{k}$ representing the idiosyncratic time-change for the $k$th price, and $\boldsymbol{\alpha}G_{0}$ representing a common time change for all prices, where $G_{0},\dots,G_{n}$ are independent gamma subordinators and $\boldsymbol{\alpha}\in[0,\infty)^{n}$, then when ${\bf T}(t)$ jumps by $\Delta{\bf T}(t)$, ${\bf Z}(t)$ jumps by the componentwise composition ${\bf B}(\Delta{\bf T}(t))$ in distribution, rather than pathwise for traditional subordination. Consequently, the WVAG process, using weak subordination, preserves the essence of traditional subordination in a distributional or weak sense.
This provides a naturally multivariate process with a flexible dependence structure where the Brownian motion has arbitrary covariance in contrast to other approaches. Specifically, these include Luciano and Semeraro LS10 who in one model use traditional subordination restricted to independent Brownian motions to remain a Lévy process, Ballotta and Bonfiglioli BB14 who use summing independent univariate VG processes $V_{0},\dots,V_{n}$ in the form ${\bf Z}=(\alpha_{1}V_{0}+V_{1},\dots,\alpha_{n}V_{0}+V_{n})$, an approach that is also found in CaFu2013 ; HZ10 , and Gardini, Sabino and Sasso gss2021 who use the above two approaches with dependent subordinators constructed by self-decomposability. The last three references include applications to spread option pricing. Our BLDP $\bf Z$}, being a sum of VG processes, includes the cases of LS10 ; BB14 but not gss2021 . A similar way of summing univariate LDOUPs to produce a bivariate energy price for spread option pricing is also considered in (benthbook, , Section 9.2.1). Our LDOUP $\bf X$} include certain restrictions of this model, although both models are more general than each other in different ways.
@@
Energy prices such as electricity are considered nontraded assets in standard arbitrage theory since electricity bought or sold on the market is instantly consumed and cannot be dynamically and frictionlessly traded to construct a replicating portfolio. Since the market is incomplete, there are infinitely many EMMs consistent with NA, so we need to postulate a model for it. Accordingly, for the pricing of energy derivatives, we use the Esscher transform gs94 to specify a parametric family of EMMs, known as the Esscher measure $\mathbb{Q}_{{\bf h}}$ indexed by the market price of risk (MPR) $\bf h$}. This is a common approach for Lévy-based models in energy markets benthbook , as well as equity markets pasc in which the incompleteness is due to the presence of jumps. A further difference between energy and equity markets is that the discounted energy price process is not required to be a $\mathbb{Q}_{{\bf h}}$-martingale, a condition in an equity market that would ordinarily determine a unique $\bf h$}, and hence an EMM $\mathbb{Q}_{{\bf h}}$. Instead, $\bf h$} must be calibrated to, for example, observed forward or call prices as discussed in Section 5 below.
@@
Here, we show how the BDLP $\bf Z$} changes under the Esscher transform into a sum of VG processes with modified parameters, which leads to the result that the class of WVAG processes is not closed under Esscher transform. In other related works, Buchmann et al. BKMS16 studies Esscher transforms for a class of strongly subordinated Brownian motions, and Michaelsen and Szimayer MiSz17 provides numerical results about pricing options on equities that follow an exponential WVAG process using the Esscher transform, although these rely on results that do not extend to energy derivatives. Our results are more general.
@@
We derive an analytical formula for the cumulant generating function (cgf) of the innovation term of a LDOUP where the BLDP is a multivariate VG process in terms of the dilogarithm function, which generalizes a result in the univariate case by Sabino sab2020 . This is a key result given in Theorem 3.3, and is surprising in light of Remark 3.6. Consequently, we obtain an analytical formula for the cgf of the log return along with conditions for this to be finite and when it can be analytically continued into the complex plane. We use this cgf to obtain a forward pricing formula, and combined with the Fourier methods of Carr and Madan CM99 and Hurd and Zhou HZ10 , it also gives pricing formulas for call and spread options, respectively.
We place our approach to spread option pricing into the context of the existing literature. In a lognormal model, pricing has been studied using closed form approximations cardur03 ; lzd10 , which started with the well-known Kirk’s approximation, and Fourier methods pelsab2014 . In the general setting using characteristic functions, closed form approximations have also been developed CaFu2013 ; pellegrino2016 , but contrasts with the Hurd-Zhou method as the latter has arbitrary precision. Our approach to pricing spread options differs from Benth and Šaltytė-Benth bsb06 where the spread is modeled directly as a univariate LDOUP as opposed to being based on (1.1). It differs from Gardini and Sabino gs22 and Gardini, Sabino and Sasso gss22b ; gss2021 where the spread is on forwards and the option is priced using a Monte Carlo method, which is more computationally expensive as opposed to being on energy and using the Hurd-Zhou method. The models are also different. The Hurd-Zhou method has also been studied for spread options on equities and interest rates in Alfeus and Schlögl AlSc2018 .
For the WVAG BDLP, we provide simulated data examples under both the true model and the fitted model. Under the true model, the spread option price is computed via the analytical formula for the cgf of the log return with the FFT method, and is compared to the Monte Carlo method, showing close agreement. We also provide an end-to-end implementation of the spread option pricing in a simulation study by fitting the model.
As we have specified parametric models for both the energy price process $\bf S$} under the real world measure $\mathbb{P}$ and the Esscher measure $\mathbb{Q}_{{\bf h}}$, it follows that our model allows the seasonality and LDOUP parameters to be estimated under $\mathbb{P}$, specifically using the MLE method by Lu Lu21 combined with the analytical formula for the cgf of the innovation term derived here, and then allows the MPR $\bf h$} to be calibrated to observed forward or call prices. Thus, we have a joint estimation and calibration method that coherently leverages both data on the $\mathbb{P}$-dynamics of the underlying and its derivative prices to fit the model. This fitting method is consistent with a large stream of literature in incomplete equity markets ek95 ; hub ; MiSz17 ; pasc ; rus15 ; sco and energy markets bd15 ; BenthPir18 ; bsb04 ; benthbook ; bs14 ; bv13 ; gzlw23 . We extend this approach to the multivariate energy setting with a numerical implementation, which is otherwise lacking in the literature. In contrast, there is another stream of literature, which we do not adopt, that fully calibrates the model under $\mathbb{Q}$ in the univariate gss22c ; MCC98 ; sab2020 ; sab2023 ; sco and multivariate BB14 ; gs22 ; gss22b ; gss2021 ; LS10 settings. While this works in the univariate case, since it effectively amounts to defining and calibrating the model under $\mathbb{Q}$ without any reference to $\mathbb{P}$, an issue arises in the multivariate case as joint parameters cannot be calibrated in this way given that multi-asset derivatives are not liquidly traded. Consequently, the joint parameters are instead estimated by calibrating to the historical log return correlation under $\mathbb{P}$, which effectively amounts to making an unjustified assumption that the $\mathbb{P}$ and $\mathbb{Q}$ correlations are equal. This is further discussed in Remark 5.3. Our method is not faced with this problem.
Then the fitted model can be used to price spread options over a panel of strikes. We find in our example that calibrating to calls is better than calibrating to forwards, and demonstrate the extent to which errors in estimating the model parameters propagate to pricing errors.
Lastly, we provide a real data example in which the model with WVAG BDLP is fit to Australian data and used to price spread options. In particular, we use electricity prices from two states, New South Wales (NSW) and Victoria, to estimate the model and the corresponding base load quarter futures prices to calibrate it, and find that the overall model has a moderately good fit based on the distribution of the innovation term and the calibration error.
@@
The paper is organized as follows. Section 2 reviews the preliminaries, including notation, definitions of the VG and WVAG processes, and useful properties of LDOUPs. Section 3 presents the main probabilistic result, giving the analytical formula for the cgf of the innovation term of a LDOUP where the BDLP is multivariate VG. Section 4 focuses on the pricing of energy derivatives and is the main part of the paper, where we specify the model, show how the BDLP changes under the Esscher transform, obtain the analytical cgf of the log return, and finally apply this to derive pricing formulas for forwards, calls, and spread options. Section 5 details the implementation details for both the simulated and real data examples, including simulation, estimation, calibration, and the pricing of spread options using the FFT method. Section 6 presents simulated data results based on both the true parameters and a simulation study using fitted parameters. Section 7 presents the real data results. Section 8 concludes with some remarks on possible extensions.
@@ 2 Preliminaries
Let $\kappa_{{\bf X}}^{\mathbb{Q}}$, ${\cal D}^{\mathbb{Q}}_{\bf X}$, $\Phi_{{\bf X}}^{\mathbb{Q}}$ and $\Psi_{{\bf X}}^{\mathbb{Q}}$ denote the respective functions and domain under $\mathbb{Q}$. Let ${\bf B}\sim BM^{n}(\boldsymbol{\mu},{\boldsymbol{\Sigma}})$ be an $n$-dimensional Brownian motion with drift $\boldsymbol{\mu}\in\mathbb{R}^{n}$ and covariance ${\boldsymbol{\Sigma}}\in\mathbb{R}^{n\times n}$. For ${\bf t}=(t_{1},\dots,t_{n})\in[0,\infty)^{n}$, define
@@ 2.1 Variance gamma and related process
A sum of VG processes (or VG convolution (VGC)) ${\bf Z}\sim{VGC}^{n}_{m}({\bf b},\boldsymbol{\mu},{\boldsymbol{\Sigma}},\boldsymbol{\eta})$ is defined by
${\bf Z}(t)=\boldsymbol{\eta}t+{\bf V}_{1}(t)+\dots+{\bf V}_{m}(t),\quad t\geq 0,$ (2.4)
@@
${\bf Z}(t)=\boldsymbol{\eta}t+{\bf V}_{0}(t)+(V_{1}(t),\dots,V_{n}(t)),\quad t\geq 0,$ (2.5)
@@
which in turn allows the process to be simulated. The marginal components of ${\bf Z}=(Z_{1},\dots,Z_{n})$ are $Z_{k}\sim VG^{1}(1/\alpha_{k},\mu_{k},\allowbreak\Sigma_{kk},\eta_{k})$, which is a general univariate VG process.
@@ 2.2 Lévy-Driven Ornstein-Uhlenbeck Processes
The solution to SDE (1.2) is
${\bf X}(t)=e^{-\lambda(t-s)}{\bf X}(s)+e^{-\lambda(t-s)}\int_{s}^{t}e^{\lambda(u-s)}\,{\rm d}{\bf Z}(\lambda u),\quad 0\leq s\leq t.$
For $0\leq s\leq t$, we follow the notations of in vstt and define the random vectors
${\bf Z}^{*}(s,t):=\int_{s}^{t}e^{\lambda(u-s)}\,{\rm d}{\bf Z}(\lambda u),\quad{\bf Z}^{*}(t):={\bf Z}^{*}(0,t).$
Note that
@@
where $b=e^{-\lambda\Delta}$, and ${\bf Z}_{b}^{(i)}\stackrel{{\scriptstyle D}}{{=}}e^{-\lambda\Delta}{\bf Z}^{*}(\Delta)$ are the iid innovations. For convenience, we instead call ${\bf Z}^{*}(\Delta)$ the innovation term, and it has a characteristic exponent given by
@@
(see (saya, , Theorem 2.2) from which (2.9) also follows). This implies that if ${\bf Z}(t)=\boldsymbol{\eta}t+{\bf Z}_{1}(t)+\dots+{\bf Z}_{m}(t)$, $t\geq 0$, where ${\bf Z}_{1},\dots,{\bf Z}_{m}$ are independent Lévy processes, then